OpenStudy (anonymous):

prove {1/n : n is a natural} U {0} is compact using the definitions

OpenStudy (anonymous):

i'll do the easy step. Let {G_alpha} be an open cover

ganeshie8 (ganeshie8):

you want to prove every open cover contains a finite subcover

OpenStudy (anonymous):

yes, im stuck on that step

OpenStudy (anonymous):

maybe contradiction?

OpenStudy (anonymous):

Suppose there is a cover that has no finite subcover

OpenStudy (zzr0ck3r):

Do you know that a closed bounded set in \(\mathbb{R}\) is compact?

OpenStudy (anonymous):

I'm not supposed to use Heine-Borel theorem. That's why it says use the definition

OpenStudy (zzr0ck3r):

oh sorry didn't read that part.

OpenStudy (zzr0ck3r):

So open cover implies finite open cover?

OpenStudy (anonymous):

finite *subcover*

OpenStudy (zzr0ck3r):

right...

ganeshie8 (ganeshie8):

contradiction looks like a nice idea

OpenStudy (zzr0ck3r):

ok so we know the sequence converges, so we can grab some N so that everything is with in some eps of 0. I think that this N could stand for our desired N. Then go by way of contradiction and assume it is not

OpenStudy (anonymous):

Let's not use sequence since this lesson is covered before that of sequence.

OpenStudy (zzr0ck3r):

Well are you ok with this? \(S=\cup_{n=1}^{\infty}\{\frac{1}{n} \mid n\in \mathbb{N}\}\)

OpenStudy (zzr0ck3r):

where \(S\) is your set.

OpenStudy (zzr0ck3r):

even with t he bad notation...

OpenStudy (anonymous):

I was gonna say S a set :)

OpenStudy (anonymous):

I was thinking it has something to do with limit point

OpenStudy (anonymous):

 It looks like that is the suggested subcover and it's finite

OpenStudy (anonymous):

:O i think that must be it!

OpenStudy (anonymous):